A certain amount of ice is supplied heat at a constant rate of 7 minutes. For the first one minute the temperatue rises uniformly with time. Then it remains constant for the next 4 minutes and again the temperature rises at uniform rate for the last two minutes. Calculate the final temperature at the end of 7 minutes.(Given L of ice = 336*10^3 J/kg and specific heat of water = 4200 J/kg-K.)

Grade:11

1 Answers

Chetan Mandayam Nayakar
12 years ago

let amount of ice be x kg, rate of supply of heat=y J/minute

let specific heat of ice be z J/kg/K, and initial temperature of ice be -w celcius

4*y=x*(L of ice)

x*z*w=y

let final temperature be T,

4200*x*T=y*2